### v1

@@ I Introduction
However, despite the notable performance of DSSMs in isolated tasks, their practical usage in real-world scenarios is likely to be hindered by limitations in memory and computing resources. This is because practical applications often involve successive arriving data, which further exacerbates the resource constraints [17]. These constraints necessitate DSSMs to exhibit continual learning (CL) capabilities. For instance, in domains like autonomous driving systems[18], where DSSMs are expected to navigate in dynamic environments, and continuously track without frequent model retraining or extensive storage of historical task data. Consequently, the development of DSSMs capable of continual learning becomes a critical step to quickly adapt to and learn from multiple tasks [19].
@@ II-B Autodifferentiable Ensemble Kalman Filters
Numerous approximation techniques exist for addressing the intractable model evidence in Eq. (3), including variational inference, MCMC sampling, Laplace, among others [32, 33, 12, 11, 13, 16, 14, 15]. This paper predominantly employs a very recent approximation approach based on the ensemble Kalman filter (EnKF) [34, 35], namely the autodifferentiable EnKF [31]. The autodifferentiable EnKF leverages the well-established EnKF for state inference while exploiting the autodifferentiation feature to simultaneously learn the parameter $\theta$. Specifically, to derive the marginal likelihood $p({\vec{\mathbf{x}}})$, the EnKF propagates $N$ equally weighted particles, denoted as ${\mathbf{z}}_{t-1}^{1:N}$, from the filtering distribution, $p_{\theta}({\mathbf{z}}_{t-1}|{\mathbf{x}}_{1:t-1})$, see Eq. (II-B), using the transition function to approximate the forecast distribution, $p_{\theta}({\mathbf{z}}_{t}|{\mathbf{x}}_{1:t-1})$:
@@ III-A Learning the First Task
We begin with the introduction to the learning of the first task, which involves classic learning and inference in DSSMs, as presented in Section II-B. Specifically, we first evaluate the log-likelihood, see Eq. (II-B) through the propagation of the emission function using the approximated forecast distribution obtained from EnKF. Then, a recognition network is introduced to infer the initial latent state from input observations [11, 8]. More concretely, we approximate the posterior distribution of the initial latent state using the introduced parametric distribution, $q_{\phi}({\mathbf{z}}_{1,0}|{\vec{\mathbf{x}}}_{1})$, where ${\mathbf{z}}_{1,0}$ denotes the initial latent state of the first task. The distribution $q_{\phi}({\mathbf{z}}_{1,0}|{\vec{\mathbf{x}}}_{1})$ is assumed to be Gaussian, with mean ($\mu_{\phi}$) and variance ($\Sigma_{\phi}$) functions modeled by a recurrent neural network (RNN)-based recognition network with parameters $\phi$. Therefore, the loss function aimed at minimizing for the first task is expressed as follows:
@@ III-B Continual Learning for Subsequent Tasks
$\!\!\!\!\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!=\!\mathcal{L}({\vec{\mathbf{x}}}_{1};\theta,\phi)\!+\!\sum_{k=1}^{j\!-\!1}\left(\!\frac{\lambda}{2}\sum_{i}M_{k,i}\!\left(\theta_{i}\!-\!\theta_{k,i}^{*}\right)^{2}\!\right),$ (9)
where $\lambda$ is a weighting hyperparameter that represents the importance of previous tasks, and $M_{k}$ is the diagonal Fisher information matrix for the $k$-th task evaluated at $\theta^{*}_{k}$, which quantifies the importance of each model parameter with respect to the corresponding task [39]. The subscript $i$ denotes the $i$-th parameter in the model.
@@
$\!\!\!\!\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!=\!\mathcal{L}({\vec{\mathbf{x}}}_{1};\theta,\phi)\!+\!\frac{\lambda}{2}\sum_{i}\left(\tilde{M}_{j\!-\!1,i}\left(\theta_{i}\!-\!\theta_{j\!-\!1,i}^{*}\right)^{2}\right).$ (10)
@@
$\!\!\!\!\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!=\!\mathcal{L}({\vec{\mathbf{x}}}_{1};\theta,\phi)\!+\!\lambda\sum_{i}\left(\Omega_{j-1,i}\left(\theta_{i}\!-\!\theta_{j\!-\!1,i}^{*}\right)^{2}\right),$ (11)
@@
$\!\!\!\!\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!=\!\mathcal{L}({\vec{\mathbf{x}}}_{1};\theta,\phi)\!+\!\lambda\sum_{i}\left(\Lambda_{j-1,i}\left(\theta_{i}\!-\!\theta_{j-1,i}^{*}\right)^{2}\right),$ (12)
@@
The regularization utilized in LwF takes a different approach from the aforementioned methods, which is often termed functional regularization rather than parameter regularization [27]. Unlike the original LwF takes a distillation loss as the regularization term for classification problems, we use the mean squared error (MSE) loss instead given the observations are continuous. Specifically, the loss function with LwF is given by:
$\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!=\!\mathcal{L}({\vec{\mathbf{x}}}_{1};\theta,\phi)\!+\!\frac{\lambda}{j\!-\!1}\sum_{k=1}^{j\!-\!1}\mathcal{L}_{\textsc{mse}}\left(\hat{{\mathbf{x}}}_{k}(\theta),\hat{{\mathbf{x}}}_{k}(\theta_{j-1}^{*})\right),$ (14)
@@ IV-A Power Consumption Dataset
To assess the proficiency of CLDSSMs in handling multiple tasks, we divide the dataset into 4 segments. Subsequently, we randomly select 1800 successive samples from each segment to form target sequences for training. Each of these sequences serves as an independent training task, employing a batch size of 32 and a sequence length of $T=50$. The remaining 200 samples in each segment are reserved as the test data for predictive analysis.
@@
(a) DSSM-first task only
(b) DSSM
(c) CLDSSM-EWC
(d) CLDSSM-MAS
(e) CLDSSM-LwF
@@ IV-B Weather Dataset
In this subsection, we conduct another evaluation using a real-world weather dataset comprising 5844 samples with four-dimensional control inputs and one-dimensional observations. This dataset captures the weather information in London from 2005 to 2020, encompassing input features such as cloud cover, sunshine conditions, global radiation, and precipitation. The observed variable is represented by temperature measured in degrees Celsius. We partition the dataset into four segments, each containing 1461 samples spanning four years and serving as an individual task. The sequence length is defined as $T=50$, and the batch size is set to 22, preserving data for approximately one year as the ground truth for predictions.
@@
(a) DSSM-first task only
(b) DSSM
(c) CLDSSM-EWC

### v2

@@ I Introduction
However, despite the notable performance of DSSMs in isolated tasks, their practical usage in real-world scenarios is likely to be hindered by limitations in memory and computing resources. This is because practical applications often involve successive arriving data, which further exacerbates the resource constraints [17]. These constraints necessitate DSSMs to exhibit continual learning (CL) capabilities. For instance, in domains like autonomous driving systems[18], where DSSMs are expected to navigate in dynamic environments, and continuously track without frequent model retraining or extensive storage of historical task data. Consequently, the development of DSSMs capable of continual learning becomes a critical step to quickly adapt to and learn from multiple tasks [19]. It is noted that, unlike multi-task learning, which jointly addresses multiple offline tasks [20], and transfer learning, which involves transferring knowledge from one task to another [21], continual learning emphasizes a model’s capacity to learn continuously from sequential data streams. For further insights, one can refer to, e.g., [22].
@@ II-B Autodifferentiable Ensemble Kalman Filters
Numerous approximation techniques exist for addressing the intractable model evidence in Eq. (3), including variational inference, MCMC sampling, Laplace, among others [35, 36, 12, 11, 13, 16, 14, 15]. This paper predominantly employs a very recent approximation approach based on the ensemble Kalman filter (EnKF) [37, 38], namely the autodifferentiable EnKF [34]. The autodifferentiable EnKF leverages the well-established EnKF for state inference while exploiting the autodifferentiation feature to simultaneously learn the parameter $\theta$. Specifically, to derive the marginal likelihood $p({\vec{\mathbf{x}}})$, the EnKF propagates $N$ equally weighted particles, denoted as ${\mathbf{z}}_{t-1}^{1:N}$, from the filtering distribution, $p_{\theta}({\mathbf{z}}_{t-1}|{\mathbf{x}}_{1:t-1})$, see Eq. (6), using the transition function to approximate the forecast distribution, $p_{\theta}({\mathbf{z}}_{t}|{\mathbf{x}}_{1:t-1})$:
@@ III-A Learning the First Task
We begin with the introduction to the learning of the first task, which involves classic learning and inference in DSSMs, as presented in Section II-B. Specifically, we first evaluate the log-likelihood, see Eq. (7) through the propagation of the emission function using the approximated forecast distribution obtained from EnKF. Then, a recognition network is introduced to infer the initial latent state from input observations [11, 8]. More concretely, we approximate the posterior distribution of the initial latent state using the introduced parametric distribution, $q_{\phi}({\mathbf{z}}_{1,0}|{\vec{\mathbf{x}}}_{1})$, where ${\mathbf{z}}_{1,0}$ denotes the initial latent state of the first task. The distribution $q_{\phi}({\mathbf{z}}_{1,0}|{\vec{\mathbf{x}}}_{1})$ is assumed to be Gaussian, with mean ($\mu_{\phi}$) and variance ($\Sigma_{\phi}$) functions modeled by a recurrent neural network (RNN)-based recognition network with parameters $\phi$. Therefore, the loss function aimed at minimizing for the first task is expressed as follows:
@@ III-B Continual Learning for Subsequent Tasks
$\!\!\!\!\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!=\!\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!+\!\sum_{k=1}^{j\!-\!1}\left(\!\frac{\lambda}{2}\sum_{i}M_{k,i}\!\left(\theta_{i}\!-\!\theta_{k,i}^{*}\right)^{2}\!\right),$ (9)
where $\lambda$ is a weighting hyperparameter that represents the importance of previous tasks, and $M_{k}$ is the diagonal Fisher information matrix for the $k$-th task evaluated at $\theta^{*}_{k}$, which quantifies the importance of each model parameter with respect to the corresponding task [42]. The parameter learned from the $k$-th task is denoted as $\theta_{k}^{*}$, and the subscript $i$ denotes the $i$-th parameter in the model.
@@
$\!\!\!\!\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!=\!\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!+\!\frac{\lambda}{2}\sum_{i}\left(\tilde{M}_{j\!-\!1,i}\left(\theta_{i}\!-\!\theta_{j\!-\!1,i}^{*}\right)^{2}\right).$ (10)
@@
$\!\!\!\!\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!=\!\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!+\!\lambda\sum_{i}\left(\Omega_{j-1,i}\left(\theta_{i}\!-\!\theta_{j\!-\!1,i}^{*}\right)^{2}\right),$ (11)
@@
$\!\!\!\!\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!=\!\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!+\!\lambda\sum_{i}\left(\Lambda_{j-1,i}\left(\theta_{i}\!-\!\theta_{j-1,i}^{*}\right)^{2}\right),$ (12)
@@
The regularization utilized in LwF takes a different approach from the aforementioned methods, which is often termed functional regularization rather than parameter regularization [30]. Unlike the original LwF, which takes a distillation loss as the regularization term for classification problems, we use the mean squared error (MSE) loss instead given the observations are continuous. Specifically, the loss function with LwF is given by:
$\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!=\!\mathcal{L}({\vec{\mathbf{x}}}_{j};\theta,\phi)\!+\!\frac{\lambda}{j\!-\!1}\sum_{k=1}^{j\!-\!1}\mathcal{L}_{\textsc{mse}}\left(\hat{{\mathbf{x}}}_{k}(\theta),\hat{{\mathbf{x}}}_{k}(\theta_{j-1}^{*})\right),$ (14)
@@ IV-A Power Consumption Dataset
To assess the proficiency of CLDSSMs in continually learning multiple tasks, we divide the dataset into four segments chronologically, corresponding to four learning sequential tasks. Subsequently, we randomly select 1800 successive samples from each segment to form 32 training target sequences with a sequence length of $T=50$. The remaining 200 samples in each segment are reserved as the test data for predictive analysis.
@@ IV-B Weather Dataset
In this subsection, we conduct another evaluation using a real-world weather dataset comprising 5844 samples with four-dimensional control inputs and one-dimensional observations. This dataset captures the weather information in London from 2005 to 2020, encompassing input features such as cloud cover, sunshine conditions, global radiation, and precipitation. The observed variable is represented by temperature measured in degrees Celsius. Similar to the approach outlined in Section IV-A, we partition the dataset into four segments, each containing 1461 samples spanning four years and serving as an individual task. The sequence length is defined as $T=50$, and the batch size is set to 22, preserving data for approximately one year as the ground truth for predictions.